INFRARED SPECTROSCOPY AND PHOTOCHEMISTRY OF NITROMETHANE ADSORBED ON ALKALI HALIDE FILMS: CONFORMERS AND QUENCHING

\underline{TONI GREEN}, ANGELA JARAMILLO, and C.A. BAUMANN, Department of Chemistry, University of Scranton, Scranton, PA 18510.

Nitromethane was adsorbed at low temperature onto sublimated alkali halide films. We have observed two conformational isomers on these surfaces, manifest as a temperature-dependent splitting of the asymmetric C-H stretching modes. The energy difference and barrier height to interconversion between these conformations were determined for molecules adsorbed on NaCl, NaBr, KCl, and KBr films over temperatures ranging from 12 K to 160 K.

Irradiation of the adsorbate at wavelengths above 250 nm does not yield any observable photoproducts, while irradiation using wavelengths below 250 nm produces a variety of photoproduct species. The surface apparently has little effect on the π - π* (198 nm gas phase) photochemistry, but quenches the n - π* photochemistry (singlet-singlet: 270 nm; singlet-triplet: 326 nm) to an undetectable level.